Find z such that 94.0% of the standard normal curve lies to the right of z. (Round your answer to two decimal places.)
z =

Find z such that 59% of the standard normal curve lies to the right of z. (Round your answer to two decimal places.)
z =
Answer #1

Solution:

Given that,

The z - distribution of the 94 % is,

1)

P( Z > z ) = 94%

1 - P( Z < z ) = 0.94

P( Z <  ) = 1 - 0.94

P( Z < z ) = 0.06

P( Z < -1.56) = 0.06

z = -1.56

2)

The z - distribution of the 59 % is,

P( Z > z ) = 59 %

1 - P( Z < z ) = 0.59

P( Z <  ) = 1 - 0.59

P( Z < z ) = 0.41

P( Z < -0.23 ) = 0.41

z = -0.23